OpenStudy (anonymous):

solve for x: square root--(4x+1)-- +1=40 dont know how to type the sign square root

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (anonymous):

\[\sqrt{4x+1}+1=40\]

OpenStudy (anonymous):

yes that is the problem!

OpenStudy (anonymous):

subtract 1, square both sides \[\sqrt{4x+1}=39\] \[4x+1=39^2=1296\] \[4x=1295\] \[x=\frac{1295}{4}\] whatever that is

OpenStudy (anonymous):

oh damn i made a mistake!

OpenStudy (anonymous):

\[39^2=1521\]

OpenStudy (anonymous):

so you get \[4x+1=1521\] \[4x=1520\] \[x=380\]

OpenStudy (anonymous):

sorry i squared 36 instead of 39

OpenStudy (anonymous):

sweet! it is correct thanks!!

OpenStudy (anonymous):

how did u ty[e the square root part?

OpenStudy (anonymous):

i am using latex to type. i think you have a symbol palette yes?